Two particles, carrying charges -q and +q and having equal masses m each, are fixed at the ends of a light rod of length L to form a dipole. The rod is clamped at an end and is placed in a uniform electric field E with the axis of the dipole along the electric field. The rod is slightly tilted and then realeased. Neglecting gravity find the time time period of small oscillations 

Grade:12

2 Answers

Kannan
9 years ago
T=2*22/7*   l/g-Eq/m
Sayali Nagaarkar
9 years ago
for a simple pendulum, T=2#(l/g)^1/2
force experienced by charge=qE
accl=F/m=qE/m , time period=2pi(mL/qE)^1/2